OpenStudy (anonymous):

given that x-y=8 and that xy<240 find the maximum value of (x+y) given that x-y=8 and that xy<240 find the maximum value of (x+y) @Mathematics

OpenStudy (anonymous):

x-y=8 x=8+y x*y<240 (y+8)*y<240 y^2+8y<240 y^2+8y-240<0 (y+20)(y-12)<0 y<12,-20 max(y)=12 max(x)=12+8=20 max(x+y)=12+20=32

OpenStudy (anonymous):

well, technically, x+y<32 if thats not <=, but ya